# Self-Supervised Learning-Based Path Planning and Obstacle Avoidance Using PPO and B-Splines in Unknown Environments

### v1

@@ Abstract
This paper introduces SmartBSP, an advanced self-supervised learning framework for real-time path planning and obstacle avoidance in autonomous robotics navigating through complex environments. The proposed system integrates Proximal Policy Optimization (PPO) with Convolutional Neural Networks (CNN) and Actor-Critic architecture to process limited LIDAR inputs and compute spatial decision-making probabilities. The robot’s perceptual field is discretized into a grid format, which the CNN analyzes to produce a spatial probability distribution. During the training process a nuanced cost function is minimized that accounts for path curvature, endpoint proximity, and obstacle avoidance. Simulations results in different scenarios validate the algorithm’s resilience and adaptability across diverse operational scenarios. Subsequently, Real-time experiments, employing the Robot Operating System (ROS), were carried out to assess the efficacy of the proposed algorithm.
@@ I Introduction
Autonomous navigation and environment mapping for autonomous vehicles has garnered considerable attention in the field of robotics and autonomous systems research [1] [2]. More specifically, path planning methodologies are divided into two primary categories: one focusing on navigation within entirely unknown environments [3], where we only have access to real-time range sensor data, and the other dedicated to environments that are either partially or completely known [4]. The interest in path planning in unknown environments is driven by the critical need for autonomous vehicles to operate effectively in situations where prior knowledge of the environment is absent or severely limited. Examples of such applications include extraterrestrial rover missions on planets like Mars, where the terrain is largely uncharted, and search-and-rescue operations in disaster-stricken areas, where rapid changes in the environment preclude the use of pre-existing maps.
@@
Additionally, the advent of Artificial Intelligence (AI) has revolutionized the way researchers tackle the path planning problem [5], [6]. By harnessing the power of AI, scholars are now able to construct neural networks that not only learn from simulated and real data but also exhibit human-like expertise in decision-making processes [7]. However, the lack of access to useful data makes training neural networks offline for path planning challenging. In such situations, self-supervised learning methods have proven to be effective [8]. Nevertheless, learning on real-time experimental data can be time-consuming and sometimes unsafe [9]. To ensure safe learning, Proximal Policy Optimization (PPO) [10] has been employed to train the path planner offline. While PPO is widely known as a method in Reinforcement Learning (RL), its application in this paper is slightly different. This study assumes that there is no transition between states (observations), meaning the impact of decisions made at any time step on future events is ignored.
@@
While the safety of the path plays a major role in the path planning task, the smoothness of the path is equally important. Due to the low maneuverability of some vehicles, paths with high curvature or sudden changes in yaw angle are not feasible. In recent years, B-splines have been employed in numerous studies to minimize the curvature of the path [11, 12, 13]. While these studies have made notable advancements in incorporating curvature and smoothness into path planning, the computational demand and possible numerical instability of their methods—especially when dealing with non-convex obstacles—present significant challenges for real-time implementation.
@@
The method proposed in this paper incrementally builds the path based on new information received from the LIDAR sensor. The main objective of the algorithm is to find the optimal positions of control points for a local B-spline that guides the vehicle toward the target while avoiding obstacles and minimizing curvature. The contributions of this paper to the field of path planning include the following:
@@
  • The proposed path planning algorithm is tailored for completely unknown environments, capable of overcoming various non-convex obstacles.
  • The method introduced here explicitly incorporates path smoothness.
@@
  • The algorithm is designed for rapid execution, enabling real-time path planning.
  • The algorithm is generalizable to different vehicles and environments by adjusting its hyperparameters, such as field of view and range.
  • The position of the target location is generalized, meaning there is no need to train the path planner for each different target location.
@@ II Related Work
In recent years, deep learning algorithms have been used in many studies to enhance traditional path planning methods. For example, in [14], Convolutional Neural Networks (CNNs) are employed to learn pre-planned paths from other offline path planning methods, marking a significant step towards leveraging image-based methodologies in path planning. In [5], Double Deep Q-Learning is utilized to navigate path planning challenges within dynamically changing environments. Sharma et al. [6] have employed Q-learning techniques to determine the optimal path in a known environment, with the specific limitation that the goal’s location is not generalized across different scenarios. In [15], Deep Q-learning and CNNs are utilized to construct paths within grid patches derived from the environment. Wang et al. [16] leveraged neural networks to improve the probability distribution of sampling in Rapidly-exploring Random Trees Star ($RRT^{*}$) algorithm, aiming to accelerate the convergence towards an optimal solution. Pflueger et al. [8] proposed an approach using inverse reinforcement learning with soft value iteration networks (SVIN) for planetary rover path planning. This method leverages deep convolutional networks and value iteration networks to handle the complexities of navigation on Mars, demonstrating its effectiveness on both grid world and realistic Mars terrain datasets. Chiang et al. [17] introduced RL-RRT, a kinodynamic motion planner that combines reinforcement learning with sampling-based planning. Their method uses deep reinforcement learning to train an obstacle-avoiding policy, which serves as a local planner, and a reachability estimator that predicts the time to reach a state in the presence of obstacles. The next section describes the methodology and steps involved in our proposed method.
@@ III Methodology
The primary objective of the proposed path planning is finding the optimal control points of B-splines, similar to [11], but diverges from the author’s previous works by integrating a novel approach that serves two purposes: 1) it accelerates the optimization process in real-time and 2) it is capable of addressing many non-convex obstacles in the environment. Central to this methodology is the consideration of both curvature and the spatial distance of the final point of each local B-spline to the target, while concurrently ensuring avoidance of detected obstacles within the sensor range.
@@
Subsequently, the circular grid undergoes a conversion process to a square grid format (Fig. 1 (b)). This conversion facilitates seamless integration with a convolutional neural network (CNN), a pivotal component of the proposed methodology. The CNN operates on the square grid input and outputs the probability distribution governing the selection of each control point for the B-spline trajectory. Notably, it is imperative to acknowledge that within this transformation framework, the angular coordinate of the circular grid corresponds to the $y$ coordinates in the square coordinate system, whereas the radial coordinate aligns with the $x$ coordinates. The rest of this section explains the main components of the proposed method.
@@ III-A Environment
In the context of an self-supervised learning problem, the environment serves as a component that processes inputs or actions alongside current states, yielding subsequent rewards or costs. Formally, we define the environment as $\mathcal{E}=(C,\mathcal{S})$, where $\mathcal{S}$ represents the observed point cloud (state) and $C$ signifies the cost function. It is noteworthy that the choice between reward maximization and cost minimization is arbitrary, as maximizing reward can be equivalently framed as minimizing cost. Additionally, it is assumed that there is no transition between states, meaning that the learning problem focuses on improving the outcome by adjusting the actions (control points of the B-spline) for each given $5\times 5$ grid configuration.
The radial coordinates of the control points are predetermined, comprising $n+1$ radial coordinates incrementally increasing from zero to the sensor range, inclusive of the origin which signifies the location of the robot at each time step. Since the location of the first and second control points is predetermined, as shown in Fig. 1 (a), the learning algorithm is assigned the responsibility of choosing among $n-1$ distinct angular coordinates (actions) for the remaining control points. The first and second control points are always at fixed positions relative to the robot to mitigate abrupt changes in the desired yaw angle. Therefore, the set of all control points (actions) $\mathcal{A}$ can be written as:
@@
Control points $\mathcal{A}_{3},...,\mathcal{A}_{n+1}$ are the geometric centers of the chosen cell grids. Given $\mathcal{A}$, the cost of the B-spline can be calculated using the following formula:
@@
In this equation, $\rho_{i}$, where $i=1,2,3$, are the weights, $C_{\text{dist}}$ represents the cost imposed due to the L2 norm between the final point of the B-spline and the final target:
@@
where $\mathcal{G}_{\text{final}}$ is the target location in $xy$-plane. The term $C_{\text{curv}}$ denotes the cost imposed by the overall curvature of the B-spline, and $C_{\text{obs}}$ denotes the cost imposed due to collision with obstacles. In this paper, we define the cost imposed by the overall curvature of B-spline as
[…]
@@
Note that all the derivatives in (4) are with respect to $\beta$. Furthermore, $C_{\text{obs}}$ is
@@ III-B Proximal Policy Optimization
Before delving into the intricacies of Proximal Policy Optimization (PPO) [10], the framework of the self-supervised learning problem can be established as follows: $(\mathcal{S},\mathcal{A},C)$ is defined where $\mathcal{S}$ denotes the current state, representing the observed circle segment captured by the range sensor; $\mathcal{A}=(\mathcal{A}_{1},...,\mathcal{A}_{n+1})$ represents the set of $n+1$ actions, i.e., control points, selected at the current state; $C$ signifies the cost associated with choosing $\mathcal{A}$ at $\mathcal{S}$, calculated via (1). Note that each $\mathcal{A}_{i}$, $i=3,...,n+1$, is sampled from a discrete distribution over $n$ possible outcomes.
PPO can be characterized as a learning method that learns by increasing the probability of selecting actions that outperform the average. The primary distinction between PPO and other reinforcement learning methods is its endeavor to restrict the steps taken in gradient descent to mitigate the risk of local optima. PPO revolutionizes policy optimization by introducing two pivotal innovations: the Clipped Surrogate Objective (CSO) and the utilization of multiple epochs of stochastic gradient ascent for each policy update. Here, the policy is defined as
@@
where $\theta$ are the parameters of the neural network and $a$ is the sampled action vector. Since each B-spline consists of $n-1$ different actions plus the first and second points, which are fixed, the probability of choosing each specific B-spline or path is as follows:
@@
This is achieved through CSO function:
@@
Note that $r(\theta)$ exceeds $1$ when the action is more probable for the current policy than for the old policy; it will range between 00 and $1$ when the action is less probable for the current policy than for the old one. The advantage function evaluates the quality of the sampled actions relative to the expected cost, which should be estimated using the critic network, as explained in the following subsection. To clarify, when the cost of the sampled path exceeds the expected cost, the advantage function is positive. Consequently, to minimize (9), the probability of that action vector should decrease in the new update. Conversely, when the cost of the sampled path is lower than the expected cost, the advantage function is negative. Thus, to minimize (9), $r_{t}(\theta)$ should increase. Furthermore, the $clip$ function constrains the update ratio $r_{t}(\theta)$ within the interval $[1-\epsilon,1+\epsilon]$, thereby preventing large steps from being taken.
@@ III-C Actor-Critic Architecture
In this paper, the authors have employed two simple CNNs to learn the policy and the expected cost at each observation, represented by a $5\times 5$ grid occupied by obstacles. Fig. 3 depicts the schematic of the architecture used as actor. The output of the actor network is a $5\times 5$ matrix. Then, each column of the output is passed through a softmax function to convert it to a probability distribution. The selected action at each column is sampled from this probability distribution. It is worth noting that the distribution obtained from the first column of the output is irrelevant since the first column corresponds to the second point of the B-spline, which is predetermined. On the other hand, the critic part takes the $5\times 5$ matrix as input and estimates the baseline, or the expected cost, to calculate the advantage function. The configurations of the critic network is represented in Fig. 4. The loss function used to update the weights of the critic network is Mean Squared Error function:
@@ III-D Target Generalization
One of the primary challenges encountered when applying CNN-based methods for path planning tasks is the incorporation of the target location into the network. Due to the limited field of view of the robot, achieving a globally optimized solution is not feasible. However, akin to how humans confront this dilemma in real-life scenarios, the robot selects a goal location within its field of view. In this approach, we calculate the distance between the final target and all five potential final points of the B-spline. Subsequently, we select the point closest to the final target, referred to as the normalized target in this study (Fig. 5). The set containing all five normalized targets is:
@@
Moreover, we train five distinct actor networks, each targeting a different potential normalized target. In other words, the $C_{\text{dist}}$ component of (1) represents the distance between the final point of the path and the normalized target. Subsequently, the network associated with the normalized target suggests a path.
@@
However, when the path to the normalized target location is obstructed, the actor fails to propose a collision-free path. In such instances, we assume that the robot has encountered either a wall or a non-convex obstacle. Although the proposed method does not offer a path for all non-convex obstacles, to address such scenarios, we shift the normalized target from the closest one to the final target to the one with the minimum cost calculated by
@@
referred to as the temporary target $\mathcal{G}_{\text{temp}}$. Here, $C_{i},i=1,\ldots,n$, represents the cost associated with normalized target $i$ as shown in Fig. 5. Note that the set of control points associated with $\mathcal{G}_{\text{temp}}$, is denoted by $\mathcal{A}_{\text{temp}}$. This adjustment will continue until the normalized target is no longer obstructed. Situations where even this adjustment of the goal fails can be categorized into two different types: (1) when all the temporary targets are obstructed, a rare occurrence, particularly when the path update rate is high or the field of view is wide, and (2) when we encounter extreme non-convexities, such as U-shaped obstacles. Here, the authors assume that none of these situations would occur.
@@
$\epsilon$ (PPO cliping constant) $0.2$

### v2

@@ Abstract
This paper introduces Smart B-Splines (SmartBSP), an advanced self-supervised learning framework for real-time path planning and obstacle avoidance in autonomous robotics navigating through complex environments. The proposed system integrates Proximal Policy Optimization (PPO) with Convolutional Neural Networks (CNN) and Actor-Critic architecture to process limited LiDAR inputs and compute spatial decision-making probabilities. The robotic vehicle’s perceptual field is discretized into a grid format, which the CNN analyzes to produce a spatial probability distribution. During the training process a nuanced cost function is minimized that accounts for path curvature, endpoint proximity, and obstacle avoidance. Simulations results in different scenarios validate the algorithm’s resilience and adaptability across diverse operational scenarios. Subsequently, Real-time experiments, employing the Robot Operating System (ROS), are carried out to assess the efficacy of the proposed algorithm. The experimental results show that the proposed algorithm is successful in different scenarios while keeping the path as smooth as possible.
@@ I Introduction
Autonomous navigation and environment mapping for autonomous vehicles has garnered considerable attention in the field of robotics and autonomous systems research [1]. More specifically, path planning methodologies are divided into two primary categories: (1) those focusing on navigation within entirely unknown environments [2], with only access to real-time range sensor data, and (2) those dedicated to environments that are either partially or completely known [3]. Classic foundations spanning grid- and sampling-based planning further contextualize these categories, including A* and heuristic search, probabilistic roadmaps, rapidly-exploring random trees and optimal variants [4, 5, 6, 7]. The interest in path planning in unknown environments is driven by the critical need for autonomous vehicles to operate effectively in situations where prior knowledge of the environment is absent or severely limited. Examples of such applications include extraterrestrial rover missions on planets like Mars, where the terrain is largely uncharted, and search-and-rescue operations in disaster-stricken areas, where rapid changes in the environment preclude the use of pre-existing maps. In such settings, classical local strategies (e.g., Bug algorithms and frontier-based exploration) have long provided baselines [8, 9].
@@
Additionally, the advent of Artificial Intelligence (AI) has revolutionized the way researchers tackle the path planning problem [10], [11]. By harnessing the power of AI, scholars are now able to construct neural networks that not only learn from simulated and real data but also exhibit human-like expertise in decision-making processes [12]. Beyond supervised pipelines, differentiable planning modules and learned planners such as Value Iteration Networks have shown promise in embedding planning structure into deep models [13]. However, the lack of access to useful data makes the offline training of neural networks challenging. In such situations, self-supervised learning methods have proven to be effective [14]. However, in situ real-time learning during field operations can be time consuming and sometimes unsafe [15]. To ensure safe learning, Proximal Policy Optimization (PPO) [16] has been employed for offline path planner training. While PPO is widely known as a method in Reinforcement Learning (RL), this study assumes that there is no transition between states (i.e., observations), that is, the impact of decisions made at any time step on future events is ignored.
@@
While the safety of the path plays a major role in the path planning task, the smoothness of the path is equally important. Due to the low maneuverability of underactuated vehicles, paths with high curvature or sudden changes in heading are not feasible. In recent years, B-splines have been employed in numerous studies to minimize the curvature of the path [17, 19, 20]. Related trajectory-optimization approaches such as CHOMP and STOMP also explicitly shape smoothness and collision costs for feasible motion generation [21, 22]. While these studies have made notable advancements in incorporating curvature and smoothness into path planning, the computational demand and possible numerical instability of their methods—especially when dealing with non-convex obstacles—present significant challenges for real-time implementation.
@@
The SmartBSP method proposed in this paper incrementally builds available path via LiDAR-based sensor feedback. The main objective of the proposed algorithm is to determine optimal positions of control points for local B-splines that guide the vehicle toward the target while avoiding obstacles and minimizing curvature. The proposed path planning algorithm is:
@@
  • tailored for completely unknown environments, capable of overcoming various non-convex obstacles.
  • designed to explicitly incorporates path smoothness.
@@
  • designed for rapid execution, enabling real-time path planning.
  • vehicle and environment-agnostic via hyperparameter adjustment (e.g., field of view and range).
  • applicable to any generic target location and, therefore, does not require path planner retraining .
@@ II Related Work
In recent years, deep learning algorithms have been used in many studies to enhance traditional path planning methods. For example, in [23], Convolutional Neural Networks (CNNs) are employed to learn pre-planned paths from other offline path planning methods, marking a significant step towards leveraging image-based methodologies in path planning. In [10], Double Deep Q-Learning is utilized to navigate path planning challenges within dynamically changing environments. Sharma et al. [11] have employed Q-learning techniques to determine the optimal path in a known environment, with the specific limitation that the goal’s location is not generalized across different scenarios. In [24], Deep Q-learning and CNNs are utilized to construct paths within grid patches derived from the environment. Wang et al. [25] leveraged neural networks to improve the probability distribution of sampling in Rapidly-exploring Random Trees Star ($RRT^{*}$) algorithm, aiming to accelerate the convergence towards an optimal solution. Pflueger et al. [14] proposed an approach using inverse reinforcement learning with soft value iteration networks (SVIN) for planetary rover path planning. This method leverages deep convolutional networks and value iteration networks to handle the complexities of navigation on Mars, demonstrating its effectiveness on both grid world and realistic Mars terrain datasets. Chiang et al. [26] introduced RL-RRT, a kinodynamic motion planner that combines reinforcement learning with sampling-based planning. Their method uses deep reinforcement learning to train an obstacle-avoiding policy, which serves as a local planner, and a reachability estimator that predicts the time to reach a state in the presence of obstacles. Complementary lines of work integrate differentiable planners (e.g., VIN) and learning–sampling hybrids (e.g., PRM-RL) for long-range navigation [13, 27]. For reactive multi-agent avoidance and local navigation baselines, velocity-obstacle methods such as ORCA are widely used [28]. On the optimization side, CHOMP and STOMP remain influential for smooth, collision-aware motion generation [21, 22]. The next section describes the methodology and steps involved in our proposed method.
@@ III Methodology
The primary objective of the proposed path planning method is to determine the optimal control points of B-splines, similar to that of [17, 18], but diverges from the authors’ previous works by integrating a novel approach that: 1) accelerates the optimization process in real-time and 2) it is capable of addressing many non-convex obstacles in an unknown environment. Central to this methodology is the consideration of both curvature and the spatial distance of the final point of each local B-spline to the target, while concurrently ensuring avoidance of detected obstacles within a given sensor range.
@@
Subsequently, the circular grid undergoes a conversion process to a square grid format (Fig. 1 (b)). This conversion facilitates seamless integration with a convolutional neural network (CNN), a pivotal component of the proposed methodology. The CNN operates on the square grid input and outputs the probability distribution governing the selection of each control point for the B-spline trajectory. Here, the angular coordinate of the circular grid corresponds to the $y$ coordinates in the square coordinate system, whereas the radial coordinate aligns with the $x$ coordinates.
@@ III-A Environment
In the context of an self-supervised learning problem, the environment serves as a component that processes inputs or actions alongside current states, yielding subsequent rewards or costs. Formally, we define the environment as $\mathcal{E}=(C,\mathcal{S})$, where $\mathcal{S}$ represents the observed point cloud (state) and $C$ signifies the cost function, noting here that the choice between reward maximization and cost minimization is arbitrary, as maximizing reward can be equivalently framed as minimizing cost. It is assumed that there is no transition between states. In other words, the learning problem focuses on improving the outcome by adjusting the actions (i.e., control points of the B-spline) for each given $5\times 5$ grid configuration.
The radial coordinates of the control points are predetermined, comprising $n+1$ radial coordinates incrementally increasing from zero to the sensor range limit, inclusive of the origin which signifies the location of the robot at each time step. Since the location of the first and second control points is predetermined, as shown in Fig. 1 (a), the learning algorithm is left to choose among n1n-1 distinct angular coordinates (actions) for the remaining control points. The first and second control points are chosen to be at fixed positions relative to the robot so as to mitigate abrupt changes in the desired heading angle. Therefore, the set of all control points (actions) $\mathcal{A}$ can be written as:
@@
Control points $\mathcal{A}_{3},...,\mathcal{A}_{n+1}$ represent the geometric centers of the chosen cell grids. Given $\mathcal{A}$, the cost of the B-spline can be calculated as
@@
In this equation, $\rho_{i}$, where $i=1,2,3$, are the weights, $C_{\text{dist}}$ represents the cost imposed due to the L2 norm between the final point of the B-spline and the final target location such that
@@
where $\mathcal{G}_{\text{final}}$ is the target location. The term $C_{\text{curv}}$ denotes the cost imposed by the overall curvature of the B-spline, and $C_{\text{obs}}$ denotes the cost imposed due to collision with obstacles. In this paper, we define the cost imposed by the overall curvature of B-spline as
$C_{\text{curv}}(\mathcal{A})=\int_{\mathcal{A}_{1}}^{\mathcal{A}_{n+1}}k(\beta)^{2}\,d\beta$ (3)
@@
noting that all the derivatives in (4) represent partial derivatives with respect to $\beta$. Furthermore, $C_{\text{obs}}$ is
@@ III-B Proximal Policy Optimization
The framework of the self-supervised learning problem is such that $(\mathcal{S},\mathcal{A},C)$ is defined where $\mathcal{S}$ denotes the current state, representing the observed circle segment captured by the range sensor; $\mathcal{A}=(\mathcal{A}_{1},...,\mathcal{A}_{n+1})$ represents the set of $n+1$ actions (i.e., control points) selected at the current state; $C$ signifies the cost associated with choosing $\mathcal{A}$ at $\mathcal{S}$, calculated via (1). Note that each $\mathcal{A}_{i}$, i=3,,n+1i=3,...,n+1, is sampled from a discrete distribution over $n$ possible outcomes.
PPO [16] can be characterized as a learning method that learns by increasing the probability of selecting actions that outperform the average. The primary distinction between PPO and other reinforcement learning methods is that it restricts the size of the steps taken in gradient descent to mitigate the risk of local optima. PPO incorporates the Clipped Surrogate Objective (CSO) and the utilization of multiple epochs of stochastic gradient ascent for each policy update. Here, the policy is defined as
@@
where $\theta$ are the parameters of the neural network and $a$ is the sampled action vector. Since each B-spline consists of n1n-1 different actions in addition to the fixed first and second points, the probability of choosing each specific B-spline or path is as follows:
@@
This is achieved through the CSO function:
@@ III-C Actor-Critic Architecture
In this paper, the authors have employed two simple CNNs to learn the policy and the expected cost at each observation, represented by a $5\times 5$ grid occupied by obstacles. Fig. 3 depicts the schematic of the architecture used as actor. The output of the actor network is a $5\times 5$ matrix. Each column of the output matrix is converted to a probability distribution via a softmax function. The selected action at each column is sampled from this probability distribution. It is noted that the distribution obtained from the first column of the output is irrelevant since the first column corresponds to the predetermined second point of the B-spline. The critic network then takes the $5\times 5$ matrix as input and estimates the baseline (i.e., the expected cost) to calculate the advantage function. The configuration of the critic network is represented in Fig. 4. The loss function used to update the weights of the critic network is Mean Squared Error function:
@@ III-D Target Generalization
One of the primary challenges encountered when applying CNN-based methods for path planning tasks is the incorporation of the target location into the network. Due to the robotic vehicle’s limited field of view, achieving a globally optimized solution is not feasible. However, akin to how humans confront this dilemma in real-life scenarios, the robot is tasked with selecting a goal location within its field of view. In this approach, the distance between the final target and all five potential final points of the B-spline are calculated. Subsequently, we select the point closest to the final target, referred to as the ”normalized target” in this study (Fig. 5). The set containing all five normalized targets is:
@@
SmartBSP trains five distinct actor networks, each targeting a different potential normalized target. In other words, the $C_{\text{dist}}$ component of (1) represents the distance between the final point of the path and the normalized target. Subsequently, the network associated with the normalized target suggests a resulting optimized path.
@@
However, when the path to the normalized target location is obstructed, the actor fails to propose a collision-free path. In such instances, it is assumed that the robot has encountered either a wall or a non-convex obstacle. The proposed method does not offer a path for all non-convex obstacles. To address such scenarios, however, the normalized target is shifted to that with the minimum cost. Cost $C_{i}(\mathcal{S},\mathcal{A})$ is calculated by
@@
This temporary target is denoted as $\mathcal{G}_{\text{temp}}$. Here, $C_{i},i=1,\ldots,n$, represents the cost associated with normalized target $i$ as shown in Fig. 5. The set of control points associated with $\mathcal{G}_{\text{temp}}$, is denoted by $\mathcal{A}_{\text{temp}}$. This adjustment (i.e., incorporating $\mathcal{G}_{\text{temp}}$ and $\mathcal{A}_{\text{temp}}$) continues until the normalized target is no longer obstructed. The SmartBSP method for obstructed paths is summarized in Algorithm 1. Situations where even this adjustment of the goal fails can be categorized into two different types: (1) when all the temporary targets are obstructed (a rare occurrence, particularly when the path update rate is high or the field of view is wide) and (2) when extreme non-convexities are preset, such as U-shaped obstacles. Here, the authors assume that neither of these situations occur.
@@
$\epsilon$ (PPO clipping constant) $0.2$